OpenStudy (anonymous):

the parametric curve has the equations x=sin2t , y=cos2t why does it move in the clockwise direction on the circle? and why does it go around twice?

zepdrix (zepdrix):

Why? :d Like you want a philosophical reasoning? :D Because ummm ... if it wasn't so.... math... wouldn't... exist... That's my way of saying, "I don't know"

zepdrix (zepdrix):

But in all seriousness, you can just plug in different t values to see the direction and stuff, ya?

zepdrix (zepdrix):

\[\large\rm t=0:\qquad (x,y)=(0,1)\]\[\large\rm t=\frac{\pi}{4}:\qquad (x,y)=(1,0)\]\[\large\rm t=\frac{\pi}{2}:\qquad (x,y)=(0,-1)\]

zepdrix (zepdrix):

Recall that in your normal unit circle, cosine generally corresponds to the x-direction, while sine corresponds to the y-direction Since it's backwards in our parameter, that must... be ... doing something.. ya

OpenStudy (anonymous):

oh right, thanks

OpenStudy (empty):

Yeah also that 2 in there tells you the frequency. So like if you go \(t=2 \pi\) normally that's one revolution for \(\sin 0 \to \sin 2 \pi\) right? But once you have that 2 in there, you only have to go for half as much time to get a full rotation, see: \[\sin(2t)\]\( t = \pi\)

OpenStudy (anonymous):

if you assume that x = cos(2t) , y = sin(2t) moves counterclockwise from (1,0) then x = sin(2t) , y = cos(2t) flips the roles of x and y, and geometrically you have the reflection about the line y = x. The reflection about the line y = x changes the direction of clockwise/counterclockwise. Also note that since x = cos(2t) , y = sin(2t) starts at (1,0) when t = 0 going ccw, x = sin(2t) , y = cos(2t) starts at (0,1) and moves cw .

OpenStudy (anonymous):

but why does it go around twice?

OpenStudy (loser66):

because domain of t is from 0 to 2pi you can see it as \(x^2 = sin^2(2t) \\y^2 = cos^2 (2t) \\x^2 +y^2 =1\) That is unit circle and domain \(0\leq t \leq 2\pi\) Put it into the table like what @zepdrix did, you can see that why it goes around twice

OpenStudy (anonymous):

The answers above are good. I will have a go at it. We are looking at x = sin(2t), and y = cos(2t) Let's analyze this expression. I am going to assume that t stands for time. To simplify matters, lets rewrite the parametric equations as x = sin(θ), y = cos (θ) But the angle θ is actually equal to 2*t, which is twice the value of time t. As time t goes from 0 to 2π each angle is double the value of time. This causes the graph to complete 2 revolutions in the time it takes for 'x = cos(t), y = sin(t)' to complete one revolution. For example When t = π, the angle is 2*π. So you have completed one revolution when t = π. When t = 2π, the angle is 4π , so you completed a second revolution when t= 2π